### v1

# A Heuristic Algorithm Based on Beam Search and Iterated Local Search for the Maritime Inventory Routing Problem

Abstract
Maritime Inventory Routing Problem (MIRP) plays a crucial role in the integration of global maritime commerce levels. However, there are still no well-established methodologies capable of efficiently solving large MIRP instances or their variants due to the high complexity of the problem. The adoption of exact methods, typically based on Mixed Integer Programming (MIP), for daily operations is nearly impractical due to the CPU time required, as planning must be executed multiple times while ensuring high-quality results within acceptable time limits. Non-MIP-based heuristics are less frequently applied due to the highly constrained nature of the problem, which makes even the construction of an effective initial solution challenging. Papageorgiou et al. (2014) introduced a single-product MIRP as the foundation for MIRPLib, aiming to provide a collection of publicly available benchmark instances. However, only a few studies that propose new methodologies have been published since then. To encourage the use of MIRPLib and facilitate result comparisons, this study presents a heuristic approach that does not rely on mathematical optimization techniques to solve a deterministic, finite-horizon, single-product MIRP. The proposed heuristic combines a variation of a Beam Search algorithm with an Iterated Local Search procedure. Among the 72 instances tested, the developed methodology can improve the best-known solution for ten instances within an acceptable CPU time.
Keywords Inventory Routing Problem  $\cdot$ Hybrid Genetic Search  $\cdot$ Metaheuristics  $\cdot$ Vendor Managed Inventory
1 Introduction
In the international shipping industry, operating companies often have to manage product transport within their fleet, respecting inventory levels in ports and vessels in an integrated way. The combination of these elements takes into account the route and scheduling of the ships with inventory management, known in the literature as the Maritime Inventory Routing Problem (MIRP) (Christiansen and Fagerholt, 2009). In this challenging combinatorial problem, the routing and scheduling steps define which ports will be visited by each vessel, in which sequence and time. Meanwhile, the inventory management step determines the inventory levels of each port and vessel at each point in time within a given planning horizon (Christiansen and Fagerholt (2009), da Costa (2018)).
Solving MIRPs is a challenging task due to its alignment with real-world complexities, such as variable demand rates, multi-product logistics, and the coordination of a central supplier with multiple consumers. Fagerholt et al. (2023) reviewed MIRP research from the past decade and proposed several future directions, including applications in energy transport, such as hydrogen, ammonia, LNG, and liquefied CO2, considering increasingly aggressive decarbonization goals. They also emphasize a key consideration when addressing this type of problem: the different planning levels, particularly between the tactical and operational levels. Strategic and tactical planning problems typically involve decisions with long-term impacts, while operational planning focuses on planning over shorter time horizons, ranging from a few days or weeks to a few months, and tends to involve much harder constraints. In addition, operational planning problems require tools capable of providing high-quality results in acceptable time frames. As a result, exact methods are generally not favored as they typically deliver optimal solutions only for small-scale instances of practical size. Another challenge in MIRP is the strong interdependence between routing decisions, where factors such as inventory levels and berth limitations trigger chain reactions throughout the supply chain.
Over the past two decades, matheuristics have emerged as a growing field in Operations Research. These are optimization algorithms that combine metaheuristics with mathematical programming techniques (Papageorgiou et al., 2018), often used in the MIRP literature. On the other hand, pure heuristics (i.e. those that do not rely on mathematical optimization procedures) are less commonly applied due to the very constrained nature of the problem, as just building an initial solution can be very challenging (Friske et al., 2022).
There is no single and well-defined definition of an MIRP, but rather numerous variants tailored to specific industrial applications (Papageorgiou et al. (2014), Andersson et al. (2010)). Given this diversity, standardizing the comparison of solution methods and establishing a taxonomy is essential. To address this issue, Papageorgiou et al. (2014) introduced the MIRPLib, a collection of publicly available benchmark single-product MIRP instances. A subset of these instances, called Group 2, serves as the focus of this study. These instances feature a planning horizon of over 120 periods, with no split pickups or deliveries. The problem setting includes a single production port, multiple consumer ports, and various classes of vessels.
In this work, our objective is to develop a pure heuristic approach for MIRP applied to instances of Group 2 by Papageorgiou et al. (2014) available in MIRPLib. The heuristic consists of a variation of the Beam Search (BS) algorithm with an Iterated Local Search (BS-ILS). Achieving this goal requires several challenging steps, including the development of a package to read MIRPLib instances and the creation of a novel solution representation.
The main contributions of this work are: (i) a comprehensive package designed to extend the applicability of MIRPLib; (ii) the introduction of an innovative solution representation for MIRP; and (iii) the development of a new approach using the BS-ILS algorithm to address a deterministic MIRP within reasonable CPU times.
This paper is organized as follows. Section 2 briefly reviews the literature on deterministic MIRP works, its variants, and heuristic applications are presented. Section 3 describes the deterministic MIRP considered in this paper. Section 4 presents the proposed BS-ILS heuristic. Section 5 reports and discusses the results of computational experiments. Finally, Section 6 concludes the work and provides directions for future research.
2 Literature Review
This section provides a concise definition of MIRP, along with an overview of key studies that use heuristic methods to address it.
2.1 Maritime Inventory Routing Problems
Ship routing plays a critical role in global maritime trade, integrating and optimizing supply chains to achieve lower costs and increased efficiency (Christiansen and Fagerholt, 2009). The MIRP combines ship routing and scheduling with inventory management. It involves determining the specific ports to be visited, the sequence of visits, and the quantities to load or unload at each port, while respecting inventory constraints and minimizing total costs (Sanghikian and Martinelli, 2020).
Real-world MIRP applications are often complex, incorporating variations such as supplier/producer or central consumer configurations, inventory constraints at specific ports, variable production/consumption rates, multiple products, and spot charters. These extensions increase computational effort, making the development of effective solution methods particularly challenging. Fagerholt et al. (2023) examined MIRP variants and extensions, addressing multiple products, finite planning horizons, speed optimization, LNG applications, uncertainty handling, and the integration of combined cargo and inventory routing.
Papageorgiou et al. (2014) introduced the MIRPLib to standardize MIRP definitions and support advanced formulations. They proposed a core model for single-product distribution, ensuring inventory constraints at all ports. Building on this foundation, additional MIRP variants were developed, including Group 2 instances — deterministic, long-horizon problems with one loading port and multiple discharge ports. This simplified model assumes that the vessels travel fully loaded to discharge regions and return empty. Detailed descriptions of the original Group 2 model are provided in Section 3.
2.2 Heuristic Approaches
Many authors have developed various methods to address this problem, including mathematical models, heuristics, metaheuristics, and matheuristics. Shaabani et al. (2023) and Fagerholt et al. (2023) provided a comprehensive review of the literature on MIRP and the algorithms used. They summarized and classified recent MIRP studies, emphasizing factors such as the size of the problem (e.g., the number of ports, vessels, products, and planning horizon) and the techniques applied to solve them. Out of the 34 studies listed by Shaabani et al. (2023) and the 53 studies listed by Fagerholt et al. (2023), 38 focus on MIRP solutions using heuristic techniques.
Ronen (2002) pioneered this area by introducing a cost-based heuristic. A decade later, Stålhane et al. (2012) proposed a multi-start local search heuristic for the LNG inventory routing problem, using greedy insertion for initialization and refining solutions via neighborhood search, branch-and-bound, or their combination. Uggen et al. (2013) implemented a fix-and-relax time decomposition heuristic for MIRPs with a single product, balancing solution quality with computational efficiency, and Goel et al. (2015) proposed an iterative heuristic for LNG ship scheduling and inventory management.
Metaheuristic approaches have become important for exploring large search spaces. Christiansen et al. (2011) integrated a construction heuristic with a Genetic Algorithm (GA) for multi-product MIRPs. Siswanto et al. (2011a) proposed a multi-heuristic methodology that achieved global optimality for smaller instances. Expanding further, Siswanto et al. (2011b) developed a multi-heuristic strategy for a ship inventory routing and scheduling problem with non-dedicated compartments. More recently, Siswanto et al. (2019) proposed a GA based on multi-heuristics for an MIRP with multiple time windows. Sheikhtajian et al. (2020) studied uncertain LNG marine IRPs and proposed a metaheuristic that combines heuristics, GA, and fuzzy modeling to improve sensitivity analysis.
Hybrid metaheuristics have also been explored. Chandra et al. (2013) combined Ant Colony Optimization with linear programming for MIPRs with optional cargoes, while Song and Furman (2013) applied a large neighborhood search (LNS) heuristic to iteratively solve subproblems. De et al. (2017) applied Particle Swarm Optimization (PSO) to sustainable MIRPs with time windows, while Zhang et al. (2018) used a Lagrangian heuristic to manage delivery time uncertainty. Al-Haidous et al. (2022) investigated a short-term LNG scheduling problem, analyzing transportation expenses and emissions using PSO.
Matheuristics, which integrate mathematical programming and heuristics, are a powerful tool for solving MIRPs. Agra et al. (2014) introduced hybrid heuristics for short-sea inventory routing, combining rolling-horizon heuristics (RHH) with local branching and feasibility pumps. Shao et al. (2015) proposed a two-stage heuristic for LNG-IRPs, integrating a RHH with MIP-based neighborhood refinement methods. Hemmati et al. (2015) and Hemmati et al. (2016) developed hybrid matheuristics for cargo and inventory routing in tramp shipping. Meanwhile, Mutlu et al. (2016) focused on the LNG Annual Delivery Program (LNG-ADP), which is a schedule of LNG deliveries for a given year, introducing a matheuristic with a unique vessel selection methodology.
Agra et al. (2016) developed a MIP-based heuristic with local search techniques, while Agra et al. (2017) proposed a matheuristic to improve solutions and ensure feasibility when the mathematical model fails. Agra et al. (2018) and Diz et al. (2019) refined MIRP decomposition algorithms using an Iterated Local Search (ILS) heuristic and robust optimization. Sanghikian and Martinelli (2020) designed a variable neighborhood search (VNS) for a multi-product MIRP, integrating LP solvers to optimize continuous variables for routing solutions.
Recent research has focused on hybrid methods, parallel strategies, and uncertainty modeling. Dong et al. (2018) developed a RHH re-optimization framework for combined maritime fleet deployment and inventory management with port-visit flexibility in roll-on/roll-off shipping. Stanzani et al. (2018) proposed a multi-start heuristic for the simultaneous pickup and delivery ship routing, incorporating two enhancement methods and a local search for further solution improvement. Furthermore, Msakni and Haouari (2018) applied an optimization-based VNS for short-term LNG delivery planning, using a greedy heuristic for initialization and iteratively solving reduced instances with a commercial solver. For LNG-ADPs problems, Li and Schütz (2020) proposed an RHH, while Li et al. (2023) incorporated transshipment and intermediate storage, using various RHH configurations.
In MIRPLib-specific studies, Papageorgiou et al. (2018) introduced several heuristics, such as RHH, $K$-opt heuristics, local branching, solution polishing and hybrid approaches, integrated with a matheuristic approach in mathematical programming to solve MIRPLib Group 2 instances. Their study resulted in new best-known solutions for 26 out of 70 instances that had not yet been proven optimal, and they established new best-known bounds for 56 instances. Munguía et al. (2019) presented the Parallel Alternating Criteria Search, an LNS-based heuristic that yielded strong results for Groups 1 and 2. Friske and Buriol (2020) implemented a multi-start heuristic with LNS for Group 1 instances, while Friske et al. (2022) employed a relax-and-fix algorithm followed by fix-and-optimize to refine MIRPLib solutions. Wang et al. (2023) introduced a bilevel parallel memetic algorithm for a maritime location inventory routing problem, tested in Group 1 instances, achieving near-best solutions with favorable deviations and faster computation times.
In contrast to the reviewed studies, which often rely on mathematical models, this study employs a heuristic approach that does not rely on mathematical optimization techniques. This method minimizes the dependency on complex decision models while maintaining practical solution quality, distinguishing itself through accessibility and applicability to real-world scenarios.
3 Problem Description
The formulation presented in this section corresponds to the MIP model for Group 2 instances available in MIRPLib (Papageorgiou, 2013). The Group 2 instances describe long-horizon deterministic single product maritime inventory routing problems. In the following description, the terms of the sets “port” ($J$) and “region” ($R$) are used as synonyms, since it is assumed that there is one port per region.
In a defined time horizon of a set $T$, a heterogeneous fleet of ships, classified into different vessel classes from set $VC$, each with distinct capacities $Q^{vc}$, is assigned to service the production ports of a set $J^{P}$ and the consumption ports of a set $J^{C}$. This operation must be done while ensuring that the minimum and maximum inventory limits in port $j\in J$ ($S_{j,t}^{min}$ and $S_{j,t}^{max}$, respectively) are respected. Each port $j\in J$ is characterized by an initial inventory $s_{j,0}$ and a specific production or consumption rate $d_{j,t}$ in period $t\in T$, along with an associated indicator parameter $\Delta_{j}$, which takes the value $+1$ for production ports and $-1$ for consumption ports. Berth limit ($B_{j}$) constraints also must be respected, which restrict the number of vessels that can attempt to load or discharge in a port $j\in J$ in a given period of time. In this MIRP, a vessel must travel at full capacity from a loading port to a discharging port, and empty from a discharging port to a loading port.
The formulation follows an arc-flow mixed-integer linear programming model, where a set of nodes $\mathcal{N}$ represents the port-time pairs. A set of arcs $A^{vc}$ associated with each vessel class $vc\in VC$ defines the possible routes between nodes. Furthermore, a set of nodes $\mathcal{N}_{s,t}$ is defined, incorporating both the source $n_{s}$ and the sink $n_{t}$ nodes. The set of inter-regional arcs $A^{vc,ir}$ represents the arcs between the production and consumption regions. For each vessel class $vc\in VC$ and for each node $n\in\mathcal{N}_{s,t}$, the incoming and outgoing arcs are denoted by $RS^{vc}_{n}$ (for reverse star) and $FS^{vc}_{n}$ (for forward star), respectively, while the set of forward star inter-regional arcs is denoted by ($FS^{vc,ir}_{n}$).
All routing costs are incorporated into $C^{vc}_{a}$ (cost for vessel of vessel class $vc\in VC$ to traverse the arc $A^{vc}$). Arc costs may vary depending on the type of arc (source, sink, waiting, inter-port). The distance $\delta_{j_{1},j_{2}}$ between ports $j_{1}$ and $j_{2}$, the cost per kilometer per vessel $C^{vc}_{km}$ according to its vessel class $vc\in VC$ and the fees $\pi_{j}$ of each port $j\in J$ are considered. Some discounts are also included: a reward for finishing the route before the time horizon ends $\rho$ and a discount $\gamma^{vc}$ for vessels of a vessel class $vc\in VC$ traveling empty from a consumption port to a production port. Inventory costs $P_{j,t}$ are included as penalties, accounting for the amount of production and consumption in each port $j\in J$ to respect inventory levels in period $t\in T$.
The continuous decision variables $\alpha_{j,t}$ account for the amount to be produced or consumed in port $j\in J$ during period $t\in T$, and $s_{j,t}$ the quantity of inventory units in port $j\in J$ at the end of period $t\in T$. The integer decision variables $x_{a}^{vc}$ represent the number of vessels in a given vessel class $vc\in VC$ that traverse a specific arc $a\in A^{vc}$.
The MIRP model (Papageorgiou, 2013) is described as follows:
$\max z=$ $\sum\limits_{vc\in VC}\sum\limits_{a\in A^{vc}}-C_{a}^{vc}x_{a}^{vc}$ (routing cost)
$+\sum\limits_{j\in J}\sum\limits_{t\in T}-P_{j,t}\alpha_{j,t}$ (inventory penalties cost) (1)
subject to
$\sum\limits_{a\in FS_{n}^{vc,ir}}x_{a}^{vc}-\sum\limits_{a\in RS_{n}^{vc,ir}}x_{a}^{vc}=\begin{cases}1&\text{if }n=n_{s},\\ -1&\text{if }n=n_{t},\\ 0&\text{if }n\in\mathcal{N}\setminus\{n_{s},n_{t}\}\end{cases}$ $n\in\mathcal{N}_{s,t},\,vc\in VC$ (2)
$s_{j,t}=s_{j,t-1}+\Delta_{j}\left(d_{j,t}-\sum\limits_{vc\in VC}\sum\limits_{a\in FS_{n}^{vc,ir}}Q^{vc}x_{a}^{vc}-\alpha_{j,t}\right)$ $n=(j,t)\in\mathcal{N}$ (3)
$S_{j,t}^{\min}\leq s_{j,t}\leq S_{j,t}^{\max}$ $n=(j,t)\in\mathcal{N}$ (4)
$\sum\limits_{vc\in VC}\sum\limits_{a\in FS_{n}^{vc,ir}}x_{a}^{vc}\leq B_{j}$ $n=(j,t)\in\mathcal{N}$ (5)
$x_{a}^{vc}\in\{0,1\}$ $v\in VC,\,a\in A^{vc,ir}$ (6)
$\alpha_{j,t}\geq 0$ $n=(j,t)\in\mathcal{N}$ (7)
$x_{a}^{vc}\in\mathbb{Z_{+}}$ $v\in VC,\,a\in A^{vc}\setminus A^{vc,ir}.$ (8)
In this model, the objective function (3) minimizes the sum of transportation costs, and penalties for lost production and stockout. Constraints (2) ensure the flow balance of vessels within each vessel class. The minimum and maximum inventory limits in ports must be respected (Constraints (3) and (4)). Constraints (5) represent the berth limit constraints, while Constraints (6) ensure that at most one vessel per vessel class travels along a particular interregional arc. Constraints (7) and (8) define the variables’ domains.
4 Proposed Heuristic
The developed heuristic is a modified version of BS (Lowerre (1976), Ow and Morton (1988), Reddy (1977)), which serves as a constructive algorithm, complemented by local search procedures that improve all the generated solutions. Once these are concluded, an ILS (Lourenço et al., 2003) is applied to possibly refine the incumbent solution.
A high-level overview of the algorithm is presented in Figure 1. The process begins with a modified version of the BS heuristic which explores a limited set of the most promising solution paths at each step. As further detailed in Section 4.2, instead of expanding all possible successors, as in a traditional breadth-first search, it maintains only a fixed number of the best candidates, reducing computational effort while directing the search towards high-quality solutions, shown as triangles. These solutions undergo a local search (LS) improvement based on a Randomized Variable Neighborhood Descent (RVND) procedure (Subramanian et al., 2010), indicated as circles. Finally, the incumbent solution is possibly further refined by the ILS approach, yielding the final solution, presented in diamond form.
Figure 1: Proposed Heuristic
4.1 Solution Representation
Solutions are represented by a single vector that stores port-vessel information, defined as a call, and the order of the vector’s calls corresponds to the chronological order of the ships’ visits to these ports. Thus, a ship requiring to service a call in a port must wait its turn in the unique list to perform the operation. Figure 2 shows an example of the solution representation, which is structured as a sequence of calls corresponding to the pair (port, vessel). The sequence is chronological, but only when it involves the same ports and vessels across the calls. For a call to occur, the preceding ones must already have taken place, with respect to the port and the vessel. In the given example, the first port serviced is port P1 by vessel V3. The second port to be serviced is also port P1 by vessel V1. This means that vessel V3 must end its operation before vessel V1 begins (considering that there is only one berth in port P1). The third call corresponds to port P3, served by vessel V2 and the fourth call corresponds to vessel V1 attending port P2. To ensure feasibility, a vessel must always travel between a production port and a consumption port, in either direction, so that there are no travels between ports of identical type.
Figure 2: Example of a solution representation Call(port, vessel).
Figure 3: Redundant solution under the single-list representation.
It should be noted that our solution representation allows a given solution to have multiple representations, leading to a redundancy, as illustrated in Figure 3. Both represented solutions correspond to the same actual solution, as the calls involve distinct ships servicing different ports. Note that it is not straightforward to check which solutions are equivalent. A specific approach has been created to address this issue with some auxiliary structures, as follows. For each element in the solution vector, four auxiliary pointers are maintained, each storing the position of the last occurrence of the vessel, the last occurrence of the port, the next occurrence of the vessel, and the next occurrence of the port. This information is preprocessed and used during local search procedures to restrict them to non-redundant operations, avoiding unnecessary computation. In addition, the time for each vessel and port is also stored in order to speed up the solution evaluation.
4.2 Beam Search Algorithm
The Harpy Speech Recognition System (Lowerre, 1976) was the first to implement what later became known as Beam Search (Ow and Morton, 1988), introducing a heuristic-based approach to efficiently explore and prune the search space in speech recognition tasks. Although the procedure was originally referred to as the Locus Model of Search, the term Beam Search was already used by Reddy (1977). BS is a method for exploring decision trees, especially when dealing with a large solution space. It involves methodically expanding a limited set of solutions simultaneously, aiming to increase the chance of finding a high quality solution with limited searching effort.
The BS algorithm, as described in Morais et al. (2024), is derived from the concept of breadth-first search. It explores a graph by expanding the most promising nodes while maintaining a fixed number of nodes at each level. The approach is structured into six stages: Initialization, Node Expansion, Successor Evaluation, Beam Selection, Iteration, and Termination and Output generation.
During the Initialization stage, the algorithm establishes the starting parameters, including the initial node, which will always be an empty list, alongside the beam width. In the Node Expansion phase, successors are formed for each active node. Each successor is a partial solution. For a given node and its corresponding solution, all valid calls that can be appended to the solution are considered. Then, for every call, a new node is created, incorporating all previous calls along with the new one. Consequently, at the $i$-$th$ level of the generated tree, every solution consists of $i$ calls. In the Successor Evaluation phase, a greedy heuristic function is employed to assess the quality of each successor. Next, the Beam Selection phase involves selecting the most promising successors based on the heuristic evaluations while adhering to the beam width limit. The process continues with the Iteration stage, repeating until no further successors can be generated from the final level within the planning horizon.
In the present study, the BS heuristic is configured to accept three distinct parameters as input: number of nodes allowed at a given level (N), maximum number of children per node (w) and the number of greedy solutions for each node (q).
The initial parameter to consider is $N$, which represents the number of nodes allowed within a specified level. The challenge lies in selecting a value that ensures an adequate array of potential initial solutions while keeping size within reasonable limits, which would otherwise result in higher CPU time. An increase in the number of ports and vessels, along with an extended instance time horizon, has been observed to lead to an increase in CPU time. For the test experiments, the $N$ used ranged according to the characteristics of each test. For the final experiments, different values of $N$ were evaluated based on the CPU time.
The second parameter, denoted as $w$, represents the upper limit on the number of children that can be generated by a single node. This parameter plays an important role in maintaining the diversity and heterogeneity among the solutions.
The third parameter, denoted as $q$, dictates the number of greedy solutions generated during the process, where one of them is deterministic, and $q-1$ is nondeterministic. This parameter directly interfaces with the greedy heuristic algorithm, which is used to assess the BS calculations. Specifically, the greedy heuristic is executed once following a deterministic approach, and subsequently, it is repeated $q-1$ times incorporating randomness into its execution to introduce variability in the solution space, as will be elaborated on in later sections.
One problem already mentioned in Section 4.1 is the fact that several identical solutions are represented in different ways. A simple procedure was devised with the aim of reducing the occurrence of solutions that share the same cost evaluations, thereby improving the efficiency of BS, as it is undesirable to occupy unnecessary space in the BS level. This procedure functions by maintaining a dictionary where the keys correspond to the scores of potential solutions. By referencing this dictionary, it is possible to determine whether a particular score has been assigned to any solution. In such cases, these solutions are excluded from further consideration. An unlikely issue might arise with the presence of two distinct solutions that share the same cost, due to the intricate nature of the problem. Nevertheless, even if this occurs, the benefits continue to outweigh the disadvantages.
4.3 The Greedy Algorithm
To effectively compare the partial solutions, i.e., the successors produced during a BS execution, a greedy heuristic is used to construct a complete solution starting from a given partial solution, as presented in Algorithm 1. The greedy heuristic algorithm is built in both deterministic and nondeterministic versions. The heuristic constructs a feasible solution using a greedy method, making each decision by selecting the most favorable immediate choice according to a defined criterion. The criterion used involved choosing the next port at risk of violating inventory levels and determining which vessel could arrive there first. Therefore, the path was designed to prioritize, in the subsequent Call, the addition of the port at risk of inventory violation and the vessel capable of reaching this port in the earliest possible time.
1
[…]
3while  $t\leq timehorizon$  do
4
5      $nextport\leftarrow$ next port to violate
6      $nextvessel\leftarrow$ vessel arrives first
7      $s_{partial}\leftarrow$ (nextport, nextvessel)
8 end while
9return $s$
Algorithm 1 Greedy Algorithm - Deterministic version
To enhance flexibility and solution exploration, a nondeterministic version of the greedy algorithm is implemented, introducing randomness at two stages of the process: the port selection and the vessel selection. This randomness mechanism allows the algorithm to explore a wider range of possible paths, helping to escape suboptimal solutions and potentially improving the quality of the final solution.
In the nondeterministic version, instead of choosing the port that first violates its capacity constraint or the ship that first serves the selected port, the inverse of violation times is used as a probabilistic criterion. This means that ports with earlier violation times and vessels with earlier arrival times have a higher probability of being selected, although the choice is not fixed. This approach enables the algorithm to allow solutions that may seem less promising in the short term but, when combined, could lead to less infeasible and higher-quality solutions.
To incorporate this strategy into BS, the heuristic is executed $q$ times. In one execution, the process is entirely deterministic, following the original selection criterion. In the remaining $q-1$ executions, randomness is applied. At the end of these $q$ executions, the median of the solutions obtained is selected to compose the final score.
Notably, throughout the BS procedure, multiple complete solutions are assessed via this greedy heuristic. To ensure that valuable solutions are retained, the decision was made to preserve the top $N$ solutions identified by the greedy heuristic, in case they surpass the ultimate solution of the BS.
4.4 Local search
The local search procedure is based on a Randomized Variable Neighborhood Descent (RVND) scheme, where each neighborhood is explored in sequence until an improvement is found. If no improvement is found, the procedure moves on to the next neighborhood, following a random order. If any neighborhood results in a better solution, the process is restarted. Furthermore, our implementation uses the first improvement strategy.
An important consideration when considering these neighborhoods is that, in some cases, certain Calls may be displaced outside the planning horizon. When this happens, these Calls are ignored during the solution evaluation as they do not contribute to the cost or feasibility within the planned period.
Six neighborhood structures were implemented and they are described below.
  • Swap: swap two Calls in the solution.
  • Relocate: a Call is relocated to a new position in the solution.
  • Replace: the port of a Call is replaced by a port of the same type.
  • Insert: A pair of Calls is inserted at the end of the solution.
  • Remove: removes a Call along with the next Call of the same ship. This procedure is carried out to preserve vessel parity. For example, consider the following sequence of Calls: (P1, V1), (P2, V3), (P3, V1), (P1, V1), (P1, V2), (P1, V3), (P3, V2). If the Call (P3, V1) is removed, it becomes necessary to eliminate another occurrence of V1 to maintain the producer-consumer balance. In this case, both (P3, V1) and (P1, V1) are removed.
  • Swap Port: the destinations of two vessels are swapped, while restricting the swap to ports of the same type.
4.5 Iterated Local Search
ILS (Lourenço et al., 2003) applies an iterative process that alternates between perturbation and local search procedures. In each iteration, a perturbation is randomly applied to the current solution, which involves making a controlled change in its structure, by altering the routing structure of the current solution, triggering a change in the cost value of the solution. The perturbation mechanism selects one the six neighborhoods defined in Section 4.4 and applies a move at random. In our implementation, such a mechanism is called twice in a row. This perturbation generates a new candidate solution which is then refined using a local search to improve its quality within the neighborhood of the current solution. The algorithm then checks if the new solution will be accepted by a pre-established Simulated Annealing (Kirkpatrick et al., 1983) criterion. Once a solution is accepted, the algorithm updates the current solution and checks for improvement. If no improvement occurs, a counter increments. When this counter exceeds the allowed number of non-improving iterations, the solution is restored, and the counter resets. This process repeats until the maximum number of iterations is reached, at which point the final solution is returned.
5 Computational Experiments
The proposed algorithm was coded in Julia (version 1.10.5) and executed on an Intel(R) Core(TM) i7-8700K CPU @ 3.70GHz running Ubuntu 20.04, and using the GNU Parallel package (Tange, 2011) to distribute one task per thread.
5.1 Instance Details
Benchmark instances and their best-known solutions are maintained on the MIRPLib website (http://mirplib.scl.gatech.edu). The metadata includes high-level information for each instance. Most of the parameters are self-explanatory and were presented previously in Section 3. To reader is referred to Papageorgiou et al. (2014) for further explanation.
A software package has been created to access and use the public instances available in MIRPLib. Upon publication of the article, this package will be published for public use.
This paper addresses Group 2 instances, representing a single-product MIRP, considering only one port per region and never involving split pickups or split deliveries.
All instances have one loading region (LR) and different number of discharging regions (DR). The instance LR1DR02VC01V6a has one load region, two discharging regions, six vessels of one vessel class available, while instance LR1DR04VC03V15b has one load region, four discharging regions, 15 vessels of three vessel classes. The final letter “a” or “b” refers to the production/demand rates in the ports. If the letter is “a” the rates are constant, and if it is “b” the rates vary during the time horizon.
According to the classification presented by Papageorgiou et al. (2018), instances are grouped into three distinct categories based on their difficulty: Easy (E), Medium (M), and Hard (H). An instance is categorized as Hard (H) if, within a time frame of 1800 seconds, no commercial MILP solver operating with default parameter settings achieves a reduction greater than 10% of the gap, as defined in the equation 9. Conversely, an instance is considered Easy (E) if at least one commercially available MILP solver can reduce the gap by more than 90% within the same time limit using the default CPLEX settings. Instances that do not meet the criteria for the Easy or Hard categories are classified as Medium (M).
The main metric used to evaluate different solution methodologies involves computing the average proportion of the relative gap that is progressively closed as time elapses. This average is calculated for all instances classified as Easy, Medium, or Hard. In this context, the relative gap is quantitatively defined as
$\min\left\{\frac{z_{\text{method}}-z_{\text{best}}}{z_{\text{best}}},1\right\},$ (9)
where $z_{method}$ refers to the value of the objective function obtained from the method currently being assessed and $z_{best}$ represents the value of the objective function corresponding to the optimal solution identified so far in the literature.
5.2 Parameter Tuning
This section outlines how the parameters of the BS approach, the greedy algorithm, and ILS were calibrated.
5.2.1 BS Parameters
To determine a promising value for the maximum number of children per node, $w$, a set of experiments was performed. Table 1 reports the average and median objective value gaps obtained from tests in which the parameter $w$ was varied from 2 to 7. These percentages reflect the aggregated gap in objective values across all runs. The best performance was achieved with $w=2$, which resulted in the lowest average and median gaps. All tests were conducted on the Easy instances, with $N=100$ and 10 seeds per instance for each value of $w$.
Table 1: Gap analysis for different $w$ values
$w$ 2 3 4 5 6 7
Average -0.70% -0.44% 1.01% 0.60% -0.21% -0.26%
Median -0.13% -0.01% 0.05% 0.00% 0.00% 0.00%
5.2.2 Greedy Heuristic Parameters
Tests were carried out to determine the value of the number of greedy solutions $q$ for each node, with the aim of maximizing performance relative to CPU time. To this end, the number of nodes allowed per level in the BS ($N$) was set to be inversely proportional to $q$, ensuring that the CPU time remained approximately constant. Figure 4 illustrates the results, showing the calibration of $q$ for the Easy instances. The score indicates the average gap obtained across 10 different seeds. The best results were achieved for $q=3$, as it yielded the smallest average gap.
Figure 4: Visual representation of the average score across instances.
5.2.3 ILS Parameters
To tune the ILS parameters, the HyperTuning.jl package (de Dios et al., 2021) was used for automated hyperparameter optimization in Julia. In order to optimize the ILS parameters, 15 instances together with five seeds were selected. A total of 120 trials were conducted and the parameter optimization process lasted for 72 hours. The parameter range, step, and results are detailed in Table 2. The results indicated the final set of parameter values that corresponds to the configuration that achieved the best overall performance.
Table 2: ILS Parameter Optimization
Parameter Definition range step result
$sa$-$in$ SA initial probability 0.75–0.99 $0.01$ $0.79$
$sa$-$fi$ SA final probability 0.01–0.10 $0.01$ $0.01$
$ils$-$\eta$ ILS number of iterations 10–1000 10 640
$ils$-$re$ ILS number of non-improving iterations before solution restore 1–20 1 4
$ils$-$pb$ ILS number of perturbations 1–5 1 2
5.3 Results Analysis
This section presents the results of the tests conducted to assess the performance of the proposed methodologies.
5.3.1 BS and Greedy Heuristic Results
Originally, this greedy heuristic served as the sole constructive approach but was unable to produce high-quality solutions, as can be observed in Table 8 (7).
As mentioned in Section 4.3, various combinations of nondeterminism were tested, applying it separately to port selection, vessel selection, or both. Thus, by incorporating a probability factor within a normal distribution and considering standard deviation, it is possible that the next port on the route might not be the first to face risk, and the vessel selected may not be the one that reaches there first. The normal distribution was used to sample values around the average, introducing variation while keeping most variability but ensuring that the majority of samples close to that remained near the central value. The configuration that achieved the best result was to apply randomness only to port selection, while maintaining deterministic vessel selection. For these tests, the 13 smallest Group 2 instances were used, with 4 different seeds, all with a 120-day horizon, and $q$ was set to 5. The size of these instances made it possible to set a large value for $N$, namely $N=1000$. This balances high-quality results with acceptable CPU time. Table 9 (7) presents the results of these tests.
Once it was established that nondeterministic methods would be used solely for the port, additional tests were performed to evaluate three distinct techniques for aggregating the results of the greedy algorithm: namely, maximum value, mean, and median. Experiments were then conducted for the 13 smallest instances using three horizon sizes. For each instance, four executions with different seeds were carried out. The number of nodes allowed per level in the BS was set to 200. This decrease is attributed to the fact that instances now take into account a time frame extending up to 360 days. The method that showed the best performance was the median. The results of these tests are shown in Table 10 (7).
Table 3 highlights the significance of the BS within the algorithm. Each column corresponds to a specific algorithm stage (After BS, After LS, and After ILS), allowing a comparison of the impact of BS on the results. The percentages indicate the improvement achieved by incorporating BS as opposed to relying solely on the greedy heuristic. Comparing the results obtained through the exclusive use of greedy heuristic versus the application of the BS algorithm, it is possible to observe that the BS approach yields an average improvement of 18.40% in the solution following the BS stage. Notably, local search has a significant impact on the results in the absence of BS, reducing the advantage of employing BS to 7.14%. When the complete heuristic is applied, the benefit of BS further decreases to 3.57% for this set of instances. However, its use remains advantageous, despite the increased computational effort.
Comparing partial MIRP solutions can be very challenging. Different evaluation criteria were tested, but incomplete solutions resulted in ships and ports being scheduled at different times, making comparisons difficult and evaluations inaccurate. Consequently, all attempts to evaluate partial solutions without completing the total solution were found to be unsatisfactory. Partial solutions were therefore completed using the greedy algorithm, enabling meaningful comparisons within the BS framework, as shown in Table 4. For a balanced comparison, both the BS and the greedy heuristic were executed for equivalent CPU time. All experiments were carried out with 15 Easy instances, $N=10$ and $10$ seeds. This approach, although computationally expensive, since it involves running a greedy algorithm for each node in the tree, demonstrated to be much more effective than other approaches for evaluating successors.
Table 3: Percentage differences between the BS with the greedy heuristic against the greedy heuristic
Instance Horizon After BS After LS After ILS
LR1_DR02_VC01_V6a $180$ $0.00$ $0.00$ $0.00$
$360$ $-17.58$ $0.13$ $0.13$
LR1_DR02_VC02_V6a $180$ $-4.92$ $-0.86$ $-0.86$
$360$ $-13.76$ $-4.72$ $-4.17$
LR1_DR02_VC03_V7a $120$ $-15.72$ $-5.44$ $-5.37$
LR1_DR02_VC03_V8a $120$ $-27.07$ $-8.52$ $-7.65$
$180$ $-25.14$ $-1.38$ $1.48$
LR1_DR02_VC04_V8a $120$ $-9.63$ $-2.73$ $-0.54$
LR1_DR02_VC05_V8a $120$ $-1.19$ $-1.58$ $-1.08$
LR1_DR03_VC03_V13b $120$ $-13.27$ $-0.95$ $-0.86$
LR1_DR03_VC03_V16a $120$ $-66.78$ $-53.68$ $-21.05$
LR1_DR04_VC03_V15a $120$ $-22.78$ $-1.41$ $0.41$
LR1_DR04_VC03_V15b $120$ $-23.91$ $-10.11$ $-5.12$
LR1_DR04_VC05_V17a $120$ $-3.91$ $-0.73$ $-0.26$
LR1_DR04_VC05_V17b $120$ $-30.25$ $-15.05$ $-8.64$
Average -18.40 -7.14 -3.57
Table 4: Comparison between the BS with and without the greedy heuristic in terms of percentage gap
Instance Horizon BS without greedy algorithm BS with greedy algorithm
After BS After LS After ILS After BS After LS After ILS
LR1_DR02_VC01_V6a 180 79.44 43.88 5.75 0.90 0.00 0.13
360 79.07 41.66 5.57 1.76 0.13 0.13
LR1_DR02_VC02_V6a 180 82.93 19.94 30.00 0.83 0.28 0.28
360 89.47 18.06 28.33 1.67 0.28 0.28
LR1_DR02_VC03_V7a 120 93.21 43.64 27.01 4.72 0.92 0.93
LR1_DR02_VC04_V8a 120 49.69 10.14 8.90 6.90 0.92 0.92
180 92.41 21.20 20.90 10.16 1.65 0.13
LR1_DR02_VC04_V8a 360 94.21 17.09 19.64 9.20 0.70 0.00
LR1_DR02_VC03_V12a 120 94.74 87.61 42.97 0.92 0.92 0.92
LR1_DR03_VC08_V13a 120 87.05 68.06 38.26 10.81 2.19 2.19
LR1_DR04_VC02_V15a 180 91.57 60.19 33.90 19.25 2.13 2.13
LR1_DR04_VC03_V15b 120 93.51 76.33 39.75 0.70 2.14 2.14
LR1_DR04_VC05_V17a 120 94.68 91.77 41.88 4.66 0.95 0.03
LR1_DR04_VC05_V17b 120 93.47 90.82 46.87 2.22 0.00 0.00
Average 89.84 54.20 21.53 5.47 2.70 2.25
5.3.2 Proposed Heuristic Results
Two studies (Papageorgiou, 2013; Papageorgiou et al., 2018) made use of MIRPLib Group 2 instances. In the first study (Papageorgiou, 2013), the authors first performed a warm-up phase to obtain an initial good solution, which was then used to warm-start the solver. Subsequently, the main experiments were run for 24 hours. In the subsequent investigation (Papageorgiou et al., 2018), the overall CPU time per scenario remains ambiguous, due to its combination of various techniques with different time periods. Papageorgiou (2013), however, emphasizes that it is difficult to find good solutions for this group in a few minutes. There is a balance between the quality of the results and the time required to obtain them. Therefore, this work aims to evaluate how heuristic works under different CPU time conditions. To accomplish this, the parameter $N$ for the BS is evaluated using three distinct values ($N$ = 10, $N$ = 100, $N$ = 1000). For each $N$, if the instance does not reach 24 hours of CPU time, it moves to the next value of $N$. If it exceeds 24 hours, it will not proceed to the next $N$ value. This approach is intended to analyze how results improve with increased CPU time using the work by Papageorgiou et al. (2018) as a baseline reference. Another important point is that the gap reported in the results below is the best gap obtained from the runs executed within 90,000 seconds. If any stage of the main algorithm was not completed within this time frame, the reported result corresponds to the last completed stage. For example, if the algorithm does not finish the ILS phase, the reported value reflects the best result achieved up to the LS, along with its respective time. Similarly, if the algorithm does not complete the LS, the reported result comes from the BS, including both the objective value and the time.
The main results for the 120, 180, and 360 periods are summarized in Tables 5, 6, and 7, respectively. The tables include key metrics such as the value of $N$, the classification of each instance, the objective values (Objval), the total costs, the percentage of the objective gap (GAP Objval) and the CPU time (in hours). These tables show the best results achieved in 90,000 seconds (approximately 25 hours). The reported results represent the best performances obtained across all the values tested for $N$. The cases indicated by (*) show results for BS and LS, as the ILS did not complete within 90,000 seconds. The proposed heuristic demonstrated improvements in the objective value of MIRPLib (Papageorgiou et al., 2018) in ten different instances when $N$ = 1000. As expected, increasing the number of nodes improves the results. However, this improvement is accompanied by a trade-off, specifically an increase in CPU time. Furthermore, regarding the classification of instances, it can be observed that those classified as “E” generally exhibit lower objective gaps compared to instances classified as “M” or “H”.
Table 5: 120-period results
Instance N Classification Objval Total Cost GAP Objval (%) Time (h)
LR1_DR02_VC01_V6a 1000 E 33809.00 33808.95 0.00 0.03
LR1_DR02_VC02_V6a 1000 E 74982.00 74986.52 0.01 0.04
LR1_DR02_VC03_V7a 1000 E 40446.00 40403.90 $-0.10$ 0.05
LR1_DR02_VC03_V8a 1000 E 43721.00 43933.34 0.48 0.04
LR1_DR02_VC04_V8a 1000 E 41657.00 41847.57 0.46 0.08
LR1_DR02_VC05_V8a 1000 E 36659.00 36616.09 $-0.12$ 0.06
LR1_DR03_VC03_V10b 1000 M 92941.00 95667.58 2.85 0.18
LR1_DR03_VC03_V13b 1000 E 124921.00 127203.29 1.79 0.43
LR1_DR03_VC03_V16a 1000 E 82837.00 91335.26 9.30 0.26
LR1_DR04_VC03_V15a 1000 E 73312.00 74714.08 1.88 0.67
LR1_DR04_VC03_V15b 1000 E 117812.00 116451.06 $-1.17$ 0.77
LR1_DR04_VC05_V17a 1000 E 72876.00 73864.03 1.34 0.89
LR1_DR04_VC05_V17b 1000 E 105766.00 103534.03 $-2.16$ 1.02
LR1_DR05_VC05_V25a 1000 E 105328.00 105995.54 0.63 2.57
LR1_DR05_VC05_V25b 1000 E 137107.00 140504.31 2.42 2.87
LR1_DR08_VC05_V38a 1000 M 166615.00 169501.81 1.70 19.83
LR1_DR08_VC05_V40a* 1000 E 178593.00 185214.23 3.57 24.94
LR1_DR08_VC05_V40b 100 M 200746.00 234070.81 14.24 4.46
LR1_DR08_VC10_V40a 100 M 185538.00 189991.43 2.34 3.43
LR1_DR08_VC10_V40b 100 M 206315.00 234282.23 11.94 4.40
LR1_DR12_VC05_V70a* 100 H 278647.00 289736.67 3.83 18.37
LR1_DR12_VC05_V70b* 100 M 308555.00 322301.84 4.27 19.98
LR1_DR12_VC10_V70a 10 H 283154.00 290673.25 2.59 12.61
LR1_DR12_VC10_V70b 10 H 295126.00 304041.16 2.93 15.71
Table 6: 180-period results
Instance N Classification Objval Total Cost GAP Objval (%) Time (h)
LR1_DR02_VC01_V6a 1000 E 52167.00 52167.21 0.00 0.07
LR1_DR02_VC02_V6a 1000 E 129372.00 129696.36 0.25 0.09
LR1_DR02_VC03_V7a 1000 M 60547.00 61950.19 2.27 0.11
LR1_DR02_VC03_V8a 1000 E 68153.00 70478.74 3.30 0.09
LR1_DR02_VC04_V8a 1000 M 66017.00 66238.44 0.33 0.17
LR1_DR02_VC05_V8a 1000 M 58619.00 58202.65 $-0.72$ 0.14
LR1_DR03_VC03_V10b 1000 M 125638.00 122230.31 $-2.79$ 0.38
LR1_DR03_VC03_V13b 1000 E 165764.00 161724.70 $-2.50$ 0.90
LR1_DR03_VC03_V16a 1000 M 143178.00 153068.94 6.46 0.62
LR1_DR04_VC03_V15a 1000 M 118621.00 119146.56 0.44 1.50
LR1_DR04_VC03_V15b 1000 E 189989.00 188776.86 $-0.64$ 1.66
LR1_DR04_VC05_V17a 1000 M 117710.00 118868.81 0.97 2.10
LR1_DR04_VC05_V17b 1000 M 159168.00 154452.82 $-3.05$ 2.27
LR1_DR05_VC05_V25a 1000 M 171620.00 172104.96 0.28 6.04
LR1_DR05_VC05_V25b 1000 M 205368.00 219522.01 6.45 6.57
LR1_DR08_VC05_V38a 100 H 274244.00 281247.89 2.49 9.00
LR1_DR08_VC05_V40a 100 H 296760.00 304072.84 2.40 9.77
LR1_DR08_VC05_V40b 100 M 337559.00 354956.21 4.90 11.61
LR1_DR08_VC10_V40a 100 H 304261.00 309323.52 1.64 9.41
LR1_DR08_VC10_V40b 100 H 331775.00 356027.44 6.81 11.99
LR1_DR12_VC05_V70a* 10 H 460566.00 506963.62 9.15 3.40
LR1_DR12_VC05_V70b* 10 H 491160.00 573240.49 14.32 3.77
LR1_DR12_VC10_V70a* 10 H 466975.00 491159.69 4.92 3.77
LR1_DR12_VC10_V70b* 10 H 470172.00 511421.68 8.07 4.18
Table 7: 360-period results
Instance N Classification Objval Total Cost GAP Objval (%) Time (h)
LR1_DR02_VC01_V6a 1000 E 108141.00 108141.00 0.00 0.31
LR1_DR02_VC02_V6a 1000 E 283031.00 283788.36 0.27 0.44
LR1_DR02_VC03_V7a 1000 H 124282.00 126940.04 2.09 0.50
LR1_DR02_VC03_V8a 1000 M 141166.00 146807.48 3.84 0.41
LR1_DR02_VC04_V8a 1000 M 138693.00 139237.84 0.39 0.84
LR1_DR02_VC05_V8a 1000 H 122598.00 122859.71 0.21 0.71
LR1_DR03_VC03_V10b 1000 H 259888.00 260807.85 0.35 1.56
LR1_DR03_VC03_V13b 1000 N 316441.00 312621.28 $-1.22$ 3.88
LR1_DR03_VC03_V16a 1000 N 327793.00 349216.16 6.13 3.56
LR1_DR04_VC03_V15a 1000 N 252710.00 256288.64 1.40 7.75
LR1_DR04_VC03_V15b 1000 N 339308.00 340710.34 0.41 7.63
LR1_DR04_VC05_V17a 1000 H 251623.00 254040.64 0.95 11.23
LR1_DR04_VC05_V17b 1000 M 304507.00 312046.69 2.42 10.18
LR1_DR05_VC05_V25a 100 H 368628.00 373871.33 1.40 8.84
LR1_DR05_VC05_V25b 100 H 410053.00 494627.59 17.10 9.79
LR1_DR08_VC05_V38a* 100 H 596969.00 683046.79 12.60 22.52
LR1_DR08_VC05_V40a* 100 H 652380.00 659967.40 1.15 20.30
LR1_DR08_VC05_V40b* 100 H 709713.00 804362.42 11.77 22.41
LR1_DR08_VC10_V40a* 100 H 663245.00 672938.28 1.44 19.92
LR1_DR08_VC10_V40b* 100 H 724513.00 794420.68 8.80 21.26
LR1_DR12_VC05_V70a* 10 H 1021389.00 1193627.62 14.43 20.07
LR1_DR12_VC05_V70b* 10 H 1093013.00 1435397.61 23.85 21.35
LR1_DR12_VC10_V70a* 10 H 1024399.00 1046678.91 2.13 19.39
LR1_DR12_VC10_V70b* 10 H 1001541.00 1306455.89 23.34 16.12
The box plots presented in Figures 5, 6, and 7 illustrate the relationship between different stages (BS, LS, and ILS) and the variable $N$ over three distinct values (10, 100, and 1000), providing relevant information on the influence of the parameter $N$ on the results. The y-axis represents the percentage of the objective gap (GAP Objval), while the x-axis denotes the varying values of $N$.
A general observation indicates that as the value of $N$ increases from 10 to 1000, there is a visible decrease in the median GAP Objval across all stages, particularly in the BS stage, which is directly affected by this parameter. This suggests that larger values of $N$ are associated with lower percentages of gaps. Furthermore, it can be concluded that while LS has a considerable influence in all the cases analyzed, whereas the effect of ILS tends to decrease as the value of $N$ increases. This observation remains consistent across various instance sizes, with the only notable variation being the increase in GAP Objval for larger instances. Finally, Additionally, the interquartile ranges for each stage illustrate that there is a significant reduction in variability as $N$ increases. At $N$ = 10, all stages exhibit a wide spread in values, with ILS showing the least variability. As $N$ increases, the number of outliers decreases, which indicates more consistent performance across stages.
Figure 5: H = 120 with different $N$ values
Figure 6: H = 180 with different $N$ values
Figure 7: H = 360 with different $N$ values
The box plots presented in Figures 8 and 9 illustrate the distribution of GapObjval and Time (h) across different values of $N$ (10, 100, 1000) and instance classifications (E, M, H). For GapObjval, instances classified as “E” consistently exhibit lower values, the median decreasing as $N$ increases. These instances are relatively easier to solve, achieving near-optimal solutions with a lower deviation. The class “M” presents higher variability, especially for smaller values of $N$, with a gradual reduction in dispersion as the size of the problem increases. In contrast, class “H” maintains a moderate but stable GapObjval distribution, with values slightly decreasing as $N$ grows, although with less visible variability compared to class “M”.
Figure 8: GAPObjval by Class and N values
Figure 9: Time by Class and N values
With respect to time (h), there is a clear increase in computational effort as $N$ grows. For Class “E”, CPU times remain low across all values of $N$, indicating efficient convergence. In Class “M”, CPU times increase significantly, especially for larger $N$ values ($N$ = 1000), where a wider spread suggests varying solver performance across different instances. The class “H” consistently requires the highest computational effort, and as $N$ grows, there is a substantial increase in CPU time, ratifying the difficulty of these instances. The CPU time distribution in Class “H” also exhibits larger variability, suggesting that some instances require substantially more computational resources.
As one of the challenges of MIRPLib Group 2 instances was achieving good results in a short time, we evaluated the GAPObjval over time to determine the results achieved after 3600 seconds, as Munguía et al. (2019) did in their work. Figures 10 to 13 show the results for easy, medium, hard and all instances.
In a first analysis, the results may indicate that Munguía’s method generally delivers better performance than ours. Below, it is presented the average performance across different instance categories and execution times (500 and 3600 seconds). Munguía’s results can be found in Munguía et al. (2019).
For easy instances, Figure 10, the gaps are smaller. At 500 seconds, Munguía et al. (2019) reached 3%, while our method had 5%. After 3600 seconds, the gaps decreased further to 2% (Munguía et al., 2019) and 4% (ours), maintaining the same relative difference.
Figure 10: GAPObjval over time and N values - Easy instances
For medium instances, Figure 11, the gaps were 10% (Munguía) and 15% (ours) at 500 seconds. At 3600 seconds, the values dropped to 5% and 11%, respectively, with Munguía et al. (2019) still showing better performance.
Figure 11: GAPObjval over time and N values - Medium instances
As expected, the hard instances yielded the highest gaps. At 500 seconds, Munguía et al. (2019) had 25%, and our method 27%. With extended execution (3600 seconds), both improved: 10% for Munguía et al. (2019) and 22% for ours.
Figure 12: GAPObjval over time and N values - Hard instances
Considering all instances, at 500 seconds, Munguía et al. (2019) achieved an average gap of 13%, while our method performed with 17%. After 3600 seconds, both methods improved significantly, but Munguía et al. (2019) maintained the advantage with 5%, compared to 13% for our method.
Figure 13: GAPObjval over time and N values - All instances
However, the observed performance gap is influenced by the computational environments in which each method was executed. While Munguía et al. (2019) ran their experiments on a 96-core Xeon X5650 cluster, our method was executed on a single thread of an i7-8700K processor. Each core of the i7-8700K delivers approximately 2.5 times the performance of a Xeon X5650 core (PassMark Software, 2024). Based on this, the total computational power available to Munguía’s method was 38 times greater than ours. Therefore, what takes us one hour to execute could be completed on their system in about 94 seconds. This means that the performance gap (GAPObjval) must be interpreted with caution. Probably, under comparable hardware conditions, our method could achieve competitive results.
6 Conclusions
The main objective of this work was to develop a heuristic approach capable of solving a MIRP without relying on mathematical optimization models, and therefore (commercial) MILP solvers, while also aiming to obtain high-quality solutions for MIRPLib Group 2 instances. To this end, we put forward a method that combines BS with ILS. The proposed BS strategy enabled the exploration of a restricted subset of promising solution paths, significantly reducing computational effort while directing the search toward high-quality solutions. This was made possible due the implementation of a newly proposed greedy algorithm, which played a crucial role in efficiently completing partial solutions during BS. Furthermore, (iterated) local search procedures demonstrated to be essential to refine and improve the quality of solutions, producing competitive results. In particular, comparative analysis with MIRPLib benchmarks revealed ten new best-known solutions.
The proposed methodology is flexible enough to easily adjust its performance by modifying the number of nodes in the BS. This allows for a controlled trade-off between smaller gaps and higher computational times, or vice versa. For smaller instances, the methodology shows strong efficiency, reaching solutions with low optimality gaps within minutes, and in some cases surpassing the quality of known benchmark results. For larger instances, maintaining a low gap requires increasing computational time (through the number of nodes in the BS). Nevertheless, satisfactory results can still be achieved in less than 24 hours.
Future work includes the development of new local search operators and acceleration mechanisms to speed up the convergence of the method. Furthermore, adaptive perturbation mechanisms could be designed to improve the diversification phase of the proposed algorithm, helping it escape local optima.
7 Comparison of Results: BS versus Greedy Heuristic
This appendix presents the comparative results between greedy heuristic and BS, along with the experimental trials conducted to determine the appropriate type of non-determinism and to select the best aggregation strategy for the greedy heuristic.
Table 8: Greedy Algorithm without BS
Instance Bound ObjVal After BS After LS After ILS GAP ObjVal (%)
LR1_DR02_VC01_V6a 180 52166.00 52167.00 52642.82 52167.21 52167.21 0.00
360 107994.00 108141.00 129438.62 108141.00 108141.00 0.00
LR1_DR02_VC02_V6a 180 128106.00 129372.00 136876.18 130603.20 130603.20 0.95
360 275347.00 283031.00 327394.47 300509.56 298807.87 5.57
LR1_DR02_VC03_V7a 120 39318.00 40446.00 49126.32 43045.13 43018.93 6.36
LR1_DR02_VC03_V8a 120 43717.00 43721.00 59676.14 47892.73 47508.72 8.66
180 66989.00 68153.00 94931.65 75464.21 73314.49 7.57
LR1_DR02_VC04_V8a 120 41277.00 41657.00 47211.03 43183.50 42206.12 1.32
LR1_DR02_VC05_V8a 120 36088.00 36659.00 38024.22 37255.92 37008.85 0.95
LR1_DR03_VC03_V13b 120 118706.00 124921.00 151117.90 129835.53 128395.38 2.78
LR1_DR03_VC03_V16a 120 71635.00 82837.00 171093.48 153981.07 116603.27 40.76
LR1_DR04_VC03_V15a 120 72108.00 73312.00 96851.16 76859.70 74896.02 2.16
LR1_DR04_VC03_V15b 120 102148.00 117812.00 153494.96 130322.79 123931.88 5.19
LR1_DR04_VC05_V17a 120 71572.00 72876.00 80193.56 74921.30 74119.90 1.71
LR1_DR04_VC05_V17b 120 84970.00 105766.00 146881.76 123865.09 116463.71 10.11
Table 9: Comparison of the performance of different types of non-determinism
Instances No Random Random Port Random Vessel Both Random
LR1_DR02_VC01_V6a 33808.95 33808.95 33808.95 33808.95
LR1_DR02_VC02_V6a 75403.72 74981.65 75403.72 75403.72
LR1_DR02_VC03_V7a 41274.88 40454.77 41142.09 40919.53
LR1_DR02_VC03_V8a 45707.88 43956.88 44391.83 44632.84
LR1_DR02_VC04_V8a 41962.49 41959.43 41969.56 41954.04
LR1_DR02_VC05_V8a 36677.10 36767.66 36718.42 36692.15
LR1_DR03_VC03_V10b 89398.66 98455.49 97672.14 98555.54
LR1_DR03_VC03_V11a 130495.12 127476.55 129629.45 128650.85
LR1_DR03_VC03_V11b 105661.44 96296.38 105158.58 100656.24
LR1_DR04_VC03_V15a 75696.28 75133.39 75451.92 75484.39
LR1_DR04_VC03_V15b 126659.43 112634.77 125544.45 121872.08
LR1_DR04_VC05_V17a 74759.18 74593.48 74897.26 74667.23
LR1_DR04_VC05_V17b 113388.55 107885.74 112136.89 111400.20
Table 10: Comparison of results using different aggregation methods.
Instance - Horizon Cost Max Cost Mean Cost Median
LR1_DR02_VC01_V6a_120 33923.34 33939.33 34266.97
LR1_DR02_VC01_V6a_180 52647.27 52481.66 52601.11
LR1_DR02_VC01_V6a_360 109414.31 110131.37 109655.42
LR1_DR02_VC02_V6a_120 77395.06 76818.15 76624.52
LR1_DR02_VC02_V6a_180 135993.47 135549.76 135545.76
LR1_DR02_VC02_V6a_360 304937.43 303904.98 302975.50
LR1_DR02_VC03_V7a_120 42910.78 42464.15 42991.15
LR1_DR02_VC03_V7a_180 67109.18 67550.13 67528.97
LR1_DR02_VC03_V7a_360 140255.62 143311.20 142994.77
LR1_DR02_VC03_V8a_120 47900.23 48310.53 48050.30
LR1_DR02_VC03_V8a_180 77256.03 78405.74 77311.96
LR1_DR02_VC03_V8a_360 174802.30 175016.07 173030.24
LR1_DR02_VC04_V8a_120 42091.77 42164.18 42218.00
LR1_DR02_VC04_V8a_180 68388.41 67871.17 67815.13
LR1_DR02_VC04_V8a_360 145597.41 145606.29 145889.97
LR1_DR02_VC05_V8a_120 37002.98 36954.36 36988.57
LR1_DR02_VC05_V8a_180 58923.26 58750.84 58762.10
LR1_DR02_VC05_V8a_360 124037.47 123887.84 123824.35
LR1_DR03_VC03_V10b_120 110858.34 111754.60 114018.19
LR1_DR03_VC03_V10b_180 145883.67 145080.40 146118.27
LR1_DR03_VC03_V10b_360 370964.74 362337.30 358205.90
LR1_DR03_VC03_V13b_120 136033.99 134909.53 135552.79
LR1_DR03_VC03_V13b_180 173509.93 172567.81 172606.80
LR1_DR03_VC03_V13b_360 341677.79 341456.29 341661.21
LR1_DR03_VC03_V16a_120 138716.00 137684.78 131597.83
LR1_DR03_VC03_V16a_180 276705.77 276930.69 274921.64
LR1_DR03_VC03_V16a_360 682519.75 682040.67 682703.66
LR1_DR04_VC03_V15a_120 82961.41 85314.85 84632.05
LR1_DR04_VC03_V15a_180 148329.62 148596.19 148713.50
LR1_DR04_VC03_V15a_360 301765.71 301804.37 301325.94
LR1_DR04_VC03_V15b_120 143300.31 143387.59 143254.13
LR1_DR04_VC03_V15b_180 212916.42 212891.02 212375.42
LR1_DR04_VC03_V15b_360 387058.14 382095.79 381325.83
LR1_DR04_VC05_V17a_120 75673.42 75783.27 75735.27
LR1_DR04_VC05_V17a_180 122075.49 121668.81 122024.03
LR1_DR04_VC05_V17a_360 261484.75 260646.48 260256.08
LR1_DR04_VC05_V17b_120 134376.88 133850.50 132773.17
LR1_DR04_VC05_V17b_180 197886.09 198980.83 199055.59
LR1_DR04_VC05_V17b_360 412700.92 408949.66 409822.29
8 Complete Final Results
This appendix provides the complete final results from the computational experiments conducted with the proposed heuristic. Detailed results of total cost and time are shown for each instance, in three different time horizons for different values of $N$.
Table 11: 120 results
Instance After BS After LS After ILS
Total Cost Time (s) Total Cost Time (s) Total Cost Time (s)
LR1_DR02_VC01_V6a
10 34.284,56 0,61 33.808,96 1,57 33.808,95 16,06
100 34.284,55 8,13 33.808,95 3,35 33.808,95 16,98
1000 34.284,55 70,62 33.808,95 23,18 33.808,95 17,87
LR1_DR02_VC02_V6a
10 78.048,16 0,77 76.887,22 1,73 76.869,91 23,41
100 75.903,99 11,38 74.981,65 3,75 74.981,65 24,69
1000 75.896,88 93,04 74.986,52 24,38 74.986,52 24,74
LR1_DR02_VC03_V7a
10 47.374,37 0,89 45.894,60 1,66 45.155,71 25,28
100 43.143,76 14,32 42.038,41 4,32 41.656,67 26,69
1000 41.229,71 113,97 40.408,26 29,97 40.403,90 27,91
LR1_DR02_VC03_V8a
10 52.117,13 0,83 49.413,12 1,63 48.163,06 19,74
100 48.431,66 12,74 46.704,93 3,86 46.169,69 21,30
1000 45.022,11 100,18 43.933,34 26,53 43.933,34 22,40
LR1_DR02_VC04_V8a
10 44.951,88 1,36 42.621,17 1,89 42.238,81 42,21
100 43.428,05 23,38 42.121,21 6,50 41.911,72 45,47
1000 42.952,19 177,68 41.906,16 51,21 41.847,57 45,45
LR1_DR02_VC05_V8a
10 38.004,70 1,14 37.204,26 1,73 37.007,02 35,53
100 37.612,54 19,02 36.754,74 5,49 36.718,29 39,49
1000 37.573,26 148,01 36.622,13 40,21 36.616,09 39,54
LR1_DR03_VC03_V10b
10 122.609,39 3,46 115.538,44 2,80 113.295,03 108,04
100 105.428,20 52,69 101.145,26 15,00 100.567,81 110,79
1000 98.539,75 418,46 95.964,03 129,69 95.667,58 111,58
LR1_DR03_VC03_V13b
10 139.702,97 11,65 134.119,65 4,98 129.738,55 224,71
100 133.767,57 119,10 129.496,89 36,25 128.214,31 233,34
1000 131.139,72 1.048,76 127.708,77 279,39 127.203,29 230,36
LR1_DR03_VC03_V16a
10 125.766,78 7,81 119.597,46 2,92 106.775,58 143,62
100 108.280,56 78,30 103.798,86 16,61 100.793,05 143,51
1000 94.871,47 659,31 92.562,11 130,70 91.335,26 141,76
LR1_DR04_VC03_V15a
10 84.461,46 19,72 77.781,46 6,36 75.525,02 296,91
100 79.069,31 172,15 75.794,80 47,38 75.019,96 302,16
1000 77.634,98 1.640,48 74.919,84 461,55 74.714,08 296,73
LR1_DR04_VC03_V15b
10 140.139,76 24,13 131.008,91 7,53 126.556,14 388,82
100 127.266,91 206,51 121.240,65 50,38 119.902,43 381,60
1000 121.172,64 1.942,67 116.840,12 465,61 116.451,06 380,13
LR1_DR04_VC05_V17a
10 78.327,31 26,56 75.297,60 8,04 74.284,04 401,14
100 77.117,73 236,15 74.464,65 68,62 74.018,68 400,30
1000 76.676,89 2.133,04 74.122,89 677,08 73.864,03 389,46
LR1_DR04_VC05_V17b
10 125.461,17 34,15 119.074,20 9,18 114.069,25 545,55
100 115.286,57 267,20 109.308,17 77,12 108.207,31 549,08
1000 108.664,72 2.534,21 104.446,49 610,35 103.534,03 533,56
LR1_DR05_VC05_V25a
10 111.685,70 82,71 107.879,79 11,16 106.129,76 904,30
100 109.926,92 725,57 106.844,71 110,32 106.208,20 941,64
1000 108.981,57 7.231,93 106.163,11 1.102,85 105.995,54 926,51
LR1_DR05_VC05_V25b
10 196.054,84 88,05 184.818,63 18,56 161.750,38 1.208,53
100 168.185,06 826,98 157.415,61 162,53 150.807,98 1.243,89
1000 148.161,95 7.797,47 142.819,71 1.367,55 140.504,31 1.184,17
LR1_DR08_VC05_V38a
10 197.239,39 505,41 181.760,20 101,33 174.617,70 5.606,91
100 177.715,18 5.068,98 171.626,89 1.112,23 170.153,78 5.755,66
1000 174.950,49 55.892,03 170.770,36 10.101,59 169.501,81 5.393,35
LR1_DR08_VC05_V40a
10 208.065,62 554,35 192.215,93 116,63 185.280,57 5.518,14
100 196.680,93 5.568,35 188.108,08 1.168,29 185.935,56 5.753,25
1000 192.555,82 79.114,92 185.214,23 10.675,62
LR1_DR08_VC05_V40b
10 293.293,36 674,77 258.690,07 182,79 245.531,78 7.189,08
100 262.252,82 6.914,43 239.513,59 1.656,55 234.070,81 7.473,46
1000 245.203,61 85.573,09
LR1_DR08_VC10_V40a
10 212.728,01 561,46 193.728,08 125,79 189.663,97 5.573,97
100 203.767,81 5.456,39 191.443,57 1.289,16 189.991,43 5.601,75
1000 196.282,68 79.164,35
LR1_DR08_VC10_V40b
10 308.628,52 648,81 267.068,36 177,62 247.671,98 7.408,65
100 263.649,25 6.566,30 238.891,00 1.600,70 234.282,23 7.671,71
LR1_DR12_VC05_V70a
10 366.907,97 4.778,66 316.282,10 993,93 299.464,95 35.033,96
100 317.174,34 56.331,11 289.736,67 9.784,83
LR1_DR12_VC05_V70b
10 420.658,15 5.162,71 347.804,46 1.097,17 337.312,30 41.955,17
100 356.860,49 61.661,63 322.301,84 10.251,69
LR1_DR12_VC10_V70a
10 340.901,25 5.319,50 295.743,85 1.254,64 290.673,25 38.812,28
100 323.554,00 58.142,12 292.180,46 12.331,28
LR1_DR12_VC10_V70b
10 400.556,07 5.433,52 315.584,26 1.427,82 304.041,16 49.687,21
100 354.136,74 63.738,30 304.932,19 13.348,48
Table 12: 180 results
Instance After BS After LS After ILS
Total Cost Time (s) Total Cost Time (s) Total Cost Time (s)
LR1_DR02_VC01_V6a
10 52.642,82 1,12 52.167,21 1,78 52.167,21 47,80
100 52.642,81 17,98 52.167,21 5,69 52.167,21 48,43
1000 52.642,80 142,47 52.167,21 44,72 52.167,21 49,39
LR1_DR02_VC02_V6a
10 136.275,02 1,51 134.692,78 1,97 133.948,87 72,30
100 131.981,58 24,47 131.060,33 7,04 131.044,55 74,20
1000 130.773,00 197,94 129.696,36 59,04 129.696,36 75,73
LR1_DR02_VC03_V7a
10 70.994,35 1,82 68.765,03 2,10 68.450,86 80,70
100 66.001,10 29,79 64.467,39 8,69 64.430,30 81,53
1000 63.256,10 236,14 61.983,46 69,20 61.950,19 77,59
LR1_DR02_VC03_V8a
10 84.017,67 1,71 80.349,55 2,07 79.625,62 63,25
100 76.436,39 26,31 74.946,37 7,97 74.695,87 65,17
1000 71.943,71 208,16 70.523,00 60,52 70.478,74 65,12
LR1_DR02_VC04_V8a
10 69.941,11 2,80 68.251,61 2,50 67.924,21 130,90
100 67.689,80 45,00 66.600,69 14,50 66.409,01 129,43
1000 67.471,99 367,49 66.245,28 131,41 66.238,44 127,61
LR1_DR02_VC05_V8a
10 59.559,55 2,41 58.792,60 2,28 58.662,78 110,56
100 59.250,49 38,72 58.498,32 11,25 58.419,33 118,47
1000 59.102,81 305,93 58.202,72 95,85 58.202,65 116,06
LR1_DR03_VC03_V10b
10 146.394,07 6,63 140.138,60 5,29 137.143,80 290,82
100 133.642,78 93,25 129.788,09 35,10 129.405,47 300,29
1000 125.562,19 764,48 122.268,37 319,07 122.230,31 281,78
LR1_DR03_VC03_V13b
10 177.821,66 22,76 170.402,34 12,50 166.383,84 628,83
100 170.358,65 205,85 166.370,12 88,70 164.996,76 629,25
1000 165.891,32 1.846,87 162.019,22 769,19 161.724,70 611,59
LR1_DR03_VC03_V16a
10 217.091,17 17,61 212.067,87 6,45 192.871,76 460,72
100 172.232,63 156,83 166.344,51 54,19 164.804,10 455,89
1000 156.169,37 1.386,48 153.462,91 397,12 153.068,94 459,00
LR1_DR04_VC03_V15a
10 137.617,16 40,00 129.925,54 15,18 125.869,05 889,36
100 124.244,15 338,43 120.451,43 125,97 119.551,59 914,99
1000 122.221,97 3.225,96 119.443,85 1.274,73 119.146,56 902,38
LR1_DR04_VC03_V15b
10 209.509,08 44,91 198.141,58 20,14 193.887,99 1.075,15
100 198.572,29 380,51 191.821,36 157,32 190.933,64 1.074,36
1000 194.218,77 3.679,88 189.204,30 1.277,53 188.776,86 1.035,38
LR1_DR04_VC05_V17a
10 124.094,13 53,77 120.584,85 20,10 119.486,38 1.295,16
100 122.440,16 451,34 119.622,72 183,41 119.131,82 1.300,36
1000 121.651,03 4.407,27 119.129,10 1.925,75 118.868,81 1.233,94
LR1_DR04_VC05_V17b
10 181.621,65 57,91 169.971,74 24,56 167.248,46 1.630,00
100 166.398,15 490,78 159.955,15 204,63 158.660,29 1.555,65
1000 159.378,47 4.811,62 154.757,12 1.790,23 154.452,82 1.566,87
LR1_DR05_VC05_V25a
10 180.832,49 150,76 175.354,46 25,40 172.672,91 3.073,58
100 177.585,91 1.312,77 173.178,18 300,10 172.353,76 3.135,68
1000 175.501,16 15.579,02 172.185,21 3.029,10 172.104,96 3.153,64
LR1_DR05_VC05_V25b
10 310.475,54 163,10 278.354,05 72,58 256.690,03 3.904,57
100 258.067,92 1.485,22 241.850,57 675,21 237.004,99 3.911,10
1000 230.590,32 14.580,05 220.912,35 5.276,42 219.522,01 3.783,77
LR1_DR08_VC05_V38a
10 374.986,84 965,00 322.878,03 413,42 308.794,22 19.560,84
100 293.240,81 9.720,98 284.625,41 3.247,97 281.247,89 19.431,67
LR1_DR08_VC05_V40a
10 351.294,39 1.018,41 328.664,56 358,74 314.715,11 20.240,06
100 322.536,79 10.382,88 307.338,73 3.608,44 304.072,84 21.178,88
LR1_DR08_VC05_V40b
10 453.192,08 1.075,42 393.778,01 618,75 377.404,28 25.046,17
100 400.234,98 11.621,11 360.090,24 5.845,84 354.956,21 24.338,27
LR1_DR08_VC10_V40a
10 366.614,62 1.013,78 341.535,32 393,07 321.955,16 19.970,46
100 321.177,49 9.996,80 311.088,00 3.848,47 309.323,52 20.026,89
LR1_DR08_VC10_V40b
10 456.494,15 1.057,38 392.869,06 631,67 373.795,14 25.637,22
100 407.400,42 11.644,42 366.700,05 5.885,47 356.027,44 25.647,93
LR1_DR12_VC05_V70a
10 581.917,48 9.023,62 506.963,62 3.224,21
LR1_DR12_VC05_V70b
10 685.889,11 9.710,07 573.240,49 3.869,13
LR1_DR12_VC10_V70a
10 547.246,29 9.727,10 491.159,69 3.852,79
LR1_DR12_VC10_V70b
10 667.435,32 10.885,22 511.421,68 4.169,28
Table 13: 360 results
Instance After BS After LS After ILS
Total Cost Time (s) Total Cost Time (s) Total Cost Time (s)
LR1_DR02_VC01_V6a
10 131.241,95 4,15 117.805,53 4,87 115.424,27 341,24
100 110.554,82 64,40 108.894,89 28,69 108.543,07 347,44
1000 109.040,08 525,90 108.141,00 260,58 108.141,00 342,42
LR1_DR02_VC02_V6a
10 302.309,98 5,68 299.318,11 6,31 298.176,81 573,27
100 292.931,08 85,87 291.667,79 42,60 291.363,17 570,37
1000 284.715,69 712,07 283.788,36 315,11 283.788,36 549,71
LR1_DR02_VC03_V7a
10 154.997,68 6,59 147.784,22 5,96 145.380,86 597,87
100 135.588,49 99,77 133.496,51 48,24 133.092,34 587,45
1000 128.179,60 834,49 126.940,04 420,70 126.940,04 551,24
LR1_DR02_VC03_V8a
10 177.638,23 5,70 171.033,03 5,73 167.193,45 449,40
100 153.335,04 87,08 149.730,10 37,00 149.505,03 454,71
1000 149.098,64 723,81 146.883,21 302,77 146.807,48 459,45
LR1_DR02_VC04_V8a 986,09
10 149.713,33 10,83 144.984,12 8,98 143.917,35 976,18
100 141.637,68 155,10 140.249,09 79,94 139.984,70 1.015,69
1000 140.392,82 1.299,51 139.266,26 766,34 139.237,84 966,39
LR1_DR02_VC05_V8a 838,75
10 126.186,66 8,93 124.510,07 7,10 123.656,41 791,69
100 124.258,36 132,86 123.178,11 67,60 123.150,39 878,25
1000 123.746,20 1.100,93 122.859,78 601,20 122.859,71 846,30
LR1_DR03_VC03_V10b
10 329.466,44 25,70 318.285,87 29,13 310.042,30 1.694,25
100 289.974,00 276,36 286.057,05 173,39 285.561,93 1.625,33
1000 264.144,66 2.525,70 261.067,44 1.539,99 260.807,85 1.563,02
LR1_DR03_VC03_V13b
10 352.033,91 71,14 334.130,85 63,17 326.090,47 3.411,63
100 326.656,96 652,04 321.280,99 555,78 319.439,22 3.727,77
1000 315.752,07 6.013,04 312.857,51 4.307,37 312.621,28 3.663,34
LR1_DR03_VC03_V16a
10 467.382,06 69,13 454.372,57 72,68 441.807,27 3.978,05
100 390.382,71 544,77 382.499,61 493,00 380.730,40 3.808,74
1000 356.252,63 5.333,36 349.810,66 3.739,82 349.216,16 3.743,24
LR1_DR04_VC03_V15a
10 305.977,18 131,15 281.978,70 93,36 264.208,59 6.613,50
100 273.724,42 1.119,42 262.976,69 1.171,03 261.189,72 6.705,07
1000 262.101,70 11.593,14 257.127,67 9.716,80 256.288,64 6.599,61
LR1_DR04_VC03_V15b
10 411.697,21 134,18 377.519,63 118,47 356.976,69 6.415,27
100 365.132,58 1.185,99 351.912,47 1.150,73 349.552,52 6.526,01
1000 347.431,99 12.248,41 341.122,95 8.663,54 340.710,34 6.540,66
LR1_DR04_VC05_V17a
10 266.056,76 181,00 257.603,78 101,31 255.334,59 9.778,30
100 258.597,06 1.550,66 255.262,39 1.243,77 254.692,16 10.599,58
1000 256.888,26 17.771,13 254.250,37 12.238,35 254.040,64 10.406,66
LR1_DR04_VC05_V17b
10 383.190,49 180,94 358.461,83 132,87 345.029,81 9.102,59
100 341.251,26 1.587,94 330.007,26 1.227,64 327.831,51 9.140,12
1000 319.529,11 18.914,28 312.507,44 8.669,68 312.046,69 9.051,42
LR1_DR05_VC05_V25a
10 402.266,74 474,72 388.250,32 146,55 375.719,26 26.062,68
100 396.689,52 4.426,28 379.579,08 1.493,77 373.871,33 25.895,11
1000 386.148,10 63.455,80
LR1_DR05_VC05_V25b
10 592.132,45 458,07 573.194,47 194,42 517.130,83 27.906,96
100 561.152,88 4.212,32 508.332,90 4.814,35 494.627,59 26.204,78
1000 497.030,22 68.865,93
LR1_DR08_VC05_V38a
10 1.087.526,69 3.385,03 1.062.096,41 1.438,03
100 841.579,75 38.991,48 683.046,79 42.092,79
LR1_DR08_VC05_V40a
10 861.616,60 4.008,59 706.365,61 7.023,42
100 694.011,53 42.493,51 659.967,40 30.600,05
LR1_DR08_VC05_V40b
10 1.030.976,88 3.270,36 895.386,29 6.193,21
100 868.382,48 37.276,41 804.362,42 43.404,17
LR1_DR08_VC10_V40a
10 897.073,76 3.600,36 746.645,15 6.872,98
100 702.624,40 42.222,05 672.938,28 29.502,56
LR1_DR08_VC10_V40b
10 1.032.161,48 3.278,19 915.942,18 5.255,66
100 867.866,13 36.219,27 794.420,68 40.328,66
LR1_DR12_VC05_V70a
10 1.486.104,09 42.160,96 1.193.627,62 30.094,65
LR1_DR12_VC05_V70b
10 1.669.098,40 35.588,94 1.435.397,61 41.286,04
LR1_DR12_VC10_V70a
10 1.148.442,90 43.714,72 1.046.678,91 26.082,87
LR1_DR12_VC10_V70b
10 1.363.261,65 35.656,00 1.306.455,89 22.382,99